Exam 2-Chapters 22 to 24 Required information The following information applies to the questions displayed below) Ross Corporation makes all sales on account. The June 30th balance sheet balance in its accounts receivable is $430,000, of which $255.000 pertain to sales that were made during June. Budgeted sales for July are $1,280,000. Ross collects 70% of sales in the month of sale, 20% in the following month, and the final 10% in the second month after the sale Part 1 of2 points What are Rosss budgeted collections for July? 20t24) Multiple Choice $1,241000 $964 500 $1,122,000
Answer #1

May receivable = 430000-255000 = 175000

June receivable = 255000*20/30 = 170000

July cash receipt = 1280000*70% = 896000

Budgeted cash receipt in july = 896000+170000+175000 = 1241000

so answer is a) $1241000